OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 –3x + 2y = 1 5 3 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

Parth (parthkohli):

You can add both equations and eliminate.

OpenStudy (anonymous):

One question at a time.

OpenStudy (anonymous):

5

Parth (parthkohli):

3y = 15 y = 15/3 = ?

Parth (parthkohli):

Exactly.

OpenStudy (anonymous):

next :)

Parth (parthkohli):

@fredloser please attempt on your own :D

OpenStudy (anonymous):

You give it a try

Parth (parthkohli):

@Romero hey I apologize for that. I didn't report you.

OpenStudy (anonymous):

first we subtract the firstt for the second?

Parth (parthkohli):

You should multiply equation 1 by 7 and equation 2 by 3. Then you may eliminate.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

please enplane :)